When your firm hires a new employee this year, it is obligated to contribute $6,125 to a defined contribution plan for that employee, one year after the hire date. The contribution must be adjusted annually for inflation. Assume that inflation will be a constant 3% a year from this point forward. What is the cost to you of hiring a 28-year old who will be with the company for 35 years if the appropriate discount rate is 8%? Please show work
Answer #1

Cost of hiring=Present value of contributions=annual contribution*(1+inflation)/(1+discount rate)*(1-((1+inflation)/(1+discount rate))^number of years)/(1-((1+inflation)/(1+discount rate)))=6125*1.03/1.08*(1-(1.03/1.08)^35)/(1-(1.03/1.08))=$1,02,162.09